A local amateur ice skater estimates that the probability she will place first in the next regional competition is 0.60. What are the odds she will win this competition?

a) 3 to 2

b) 7 to 3

c) 3 to 8

d) 2 to 3

e) 8 to 3

f) None of the above.
Answer #1

Solution :

p = 0.60

The odds she will win this competition = p / (1 + p) = 0.60 / (1 + 0.60) = 0.60 / 1.60 = 3 / 8 = 3 to 8

C)